A jet of water squirts out horizontally from a
hole on the side of the tank as shown below.
If the hole has a diameter of 3.75 mm , what
is the height of the water above the hole in
the tank?
Answer in units of cm.
1
Expert's answer
2020-02-07T10:24:03-0500


Assume that y = 1.23 m and x = 0.546 m.


"t^2=\frac{2y}{g}=\frac{2(1.23)}{9.8}"

"t=0.501\ s"

"v=\frac{x}{t}=\frac{0.546}{0.501}=1.09\frac{m}{s}"

We have:


"0.5\rho v^2=\rho gh"

"h=\frac{v^2}{2g}=\frac{1.09^2}{2(9.8)}=0.0606\ m =6.06\ cm"